# Inside the black box: Neural network-based real-time prediction of US recessions

### v2

@@ 6 Conclusion
This research focuses on examining the real-time predictability of neural network models, compared to linear models, with respect to the recent two US recessions, the Great Recession and the Covid-19 recession. Three different neural network models are trained and updated throughout the study: a standard feed-forward neural network model (FFN), as well as two types of recurrent neural networks, LSTM and GRU, designed to effectively capture temporal dependencies in time series data. The performance of these models is evaluated using out-of-sample forecasts and compared to standard and ridge logit models. Additionally, the SHAP method is utilized to rank the predictors based on their importance for each forecast horizon, providing initial insights into the most influential predictors. The results are then compared to SHAP results obtained from the ridge logit model. The main finding are validated using both the LIME approach and the marginal effect method. For in-depth analysis, the most influential variable for the long-term forecast horizon, which is the term spread, is chosen to investigate the impact of the variable on the recession probability.
@@
This paper introduces two main contributions. Firstly, it focuses on LSTM and GRU, specialized recurrent neural network models that address issues like exploding and vanishing gradients in standard recurrent neural networks. The performance of these models is compared to FFN and to the logit variants in the context of recession forecasting. Secondly, the paper employs the SHAP method to assess the variable importance in GRU and the ridge logit model for different forecast horizons. Given the commonly perceived black box nature of neural networks, the SHAP method delivers a theoretically sound framework to provide insights into the underlying rationales influencing the predictions. The three main findings are as follows: Firstly, LSTM and GRU demonstrate strong out-of-sample performance in recession forecasting across five forecast horizons, particularly excelling in long-term predictions. Secondly, there are differences in how GRU and the ridge logit model evaluate the variable importance. The variable order differs for each forecast horizon between GRU and the ridge model. While some variables are significant in both models, their specific ranking varies notably. This suggests that GRU and the ridge model assign different weights to variables, potentially due to the neural network’s capacity to capture nonlinear relationships and assign varied weights based on distinct patterns. Furthermore, the SHAP values of GRU models show a more balanced distribution among variables compared to the ridge model. This is evident in the smaller variation in medians between variables. The difference is particularly noticeable in the short-term scenario (up to 3 months), where the ridge model heavily relies on the S&P 500 index. Moreover, the ridge model displays numerous outliers in SHAP values, indicating high sensitivity to changes in macroeconomic and financial conditions. In contrast, this emphasizes GRU’s robustness. The within variation of GRU-based SHAP values for some variables is significantly larger than for others. Certain variables show minimal change in SHAP values, while others exhibit considerable variation. This contrast is less pronounced in the ridge model outcomes, where variances are more evenly distributed, reflected in similar box sizes across variables. This suggests that the GRU has a more adaptable modeling framework, capable of adjusting variable weights in response to shifts in economic conditions. Conversely, the ridge model has relatively fixed variable orders, with limited flexibility for variation. Lastly, although the primary predictors for GRU and ridge logit models show slight differences, key indicators such as the S&P 500 index, real GDP, and private residential fixed investment consistently play a significant role in short-term predictions (up to 3 months). For longer-term forecasts (6 months or more), the term spread and producer price index become more prominent. These results are supported by other interpretation methods, the local interpretable model-agnostic explanations (LIME) for GRU and the marginal effects for the ridge logit model.

### v3

@@ 6 Conclusion
This research examines the real-time predictability of neural network models, compared to linear models, for the recent two US recessions, the Great Recession and the COVID-19 recession. Three different neural network models are trained and updated throughout the study: a standard feed-forward neural network model (FFN), as well as two types of recurrent neural networks, LSTM and GRU, designed to capture temporal dependencies in time series data effectively. The performance of these models is evaluated using out-of-sample forecasts and compared to standard and ridge logit models. Additionally, the SHAP method ranks the predictors based on their importance for each forecast horizon, providing initial insights into the most influential predictors. The results are then compared to SHAP results obtained from the ridge logit model. The main findings are validated using the LIME and marginal effect methods. For in-depth analysis, the term spread, the most influential variable for the long-term forecast horizon, is chosen to investigate the impact of the variable on the recession probability.
This paper introduces two main contributions. Firstly, it focuses on LSTM and GRU, specialized recurrent neural network models that address issues like exploding and vanishing gradients in standard recurrent neural networks. The performance of these models is compared to FFN and the logit variants in the context of recession forecasting. Secondly, the paper employs the SHAP method to assess the variable importance in GRU and the ridge logit model for different forecast horizons. Given neural networks’ commonly perceived black-box nature, the SHAP method delivers a theoretically sound framework to provide insights into the underlying rationales influencing the predictions.
@@
The three main findings are as follows: Firstly, LSTM and GRU demonstrate strong out-of-sample performance in recession forecasting across five forecast horizons, particularly in long-term predictions. Secondly, there are differences in how GRU and the ridge logit model evaluate the variable importance. The variable order differs between GRU and the ridge model for each forecast horizon. While some variables are significant in both models, their ranking varies notably. This suggests that GRU and the ridge model assign different weights to variables, potentially due to the neural network’s capacity to capture nonlinear relationships and assign varied weights based on distinct patterns. Furthermore, the SHAP values of GRU models show a more balanced distribution among variables than the ridge model. This is evident in the smaller variation in medians between variables. The difference is particularly noticeable in the short-term scenario (up to 3 months), where the ridge model heavily relies on the S&P 500 index. Moreover, the ridge model displays numerous outliers in SHAP values, indicating high sensitivity to changes in macroeconomic and financial conditions. In contrast, this emphasizes GRU’s robustness. The within variation of GRU-based SHAP values for some variables is significantly larger than for others. Certain variables show minimal change in SHAP values, while others exhibit considerable variation. This contrast is less pronounced in the ridge model outcomes, where variances are more evenly distributed, reflected in similar box sizes across variables. This suggests that the GRU has a more adaptable modeling framework capable of adjusting variable weights in response to shifts in economic conditions. Conversely, the ridge model has relatively fixed variable orders with limited flexibility for variation. Lastly, although the primary predictors for GRU and ridge logit models show slight differences, key indicators such as the S&P 500 index, real GDP, and private residential fixed investment consistently play a significant role in short-term predictions (up to 3 months). For longer-term forecasts (6 months or more), the term spread and producer price index become more prominent. Other interpretation methods support these results, such as the local interpretable model-agnostic explanations (LIME) for GRU and the marginal effects for the ridge logit model.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used Grammarly in order to improve language and readability. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.